### v1

# Graph Transformers Dream of Electric Flow

Abstract
We show theoretically and empirically that the linear Transformer, when applied to graph data, can implement algorithms that solve canonical problems such as electric flow and eigenvector decomposition. The input to the Transformer is simply the graph incidence matrix; no other explicit positional encoding information is provided. We present explicit weight configurations for implementing each such graph algorithm, and we bound the errors of the constructed Transformers by the errors of the underlying algorithms. Our theoretical findings are corroborated by experiments on synthetic data. Additionally, on a real-world molecular regression task, we observe that the linear Transformer is capable of learning a more effective positional encoding than the default one based on Laplacian eigenvectors. Our work is an initial step towards elucidating the inner-workings of the Transformer for graph data.
1 Introduction
The Transformer architecture (Vaswani et al., 2017) has seen great success in the design of graph neural networks (GNNs) (Dwivedi and Bresson, 2020; Ying et al., 2021; Kim et al., 2022; Muller et al., 2023), and it has demonstrated impressive performance in many applications ranging from prediction of chemical properties to analyzing social networks (Yun et al., 2019; Dwivedi et al., 2023; Rong et al., 2020). In contrast to this empirical success, the underlying reasons for why Transformers adapt well to graph problems are less understood. Several works study the expressivity of attention-based architectures (Kreuzer et al., 2021; Kim et al., 2021, 2022; Ma et al., 2023), but such analyses often rely on the ability of neural networks to approximate arbitrary functions, and may require a prohibitive number of parameters.
This paper is motivated by the need to understand, at a mechanistic level, how the Transformer processes graph-structured data. Specifically, we study the ability of a linear Transformer to solve certain classes of graph problems. The linear Transformer is similar to the standard Transformer, but softmax-based activation is replaced with linear attention, and the MLP layers are replaced by linear layers. Notably, the linear Transformer contains no a priori knowledge of the graph structure; all information about the graph is provided via an incidence matrix $B$. For unweighted graphs, the columns of $B$ are just $\left\{-1,0,1\right\}$-valued indicator vectors that encode whether an edge touches a vertex; no other explicit positional or structural encodings are provided.
Even in this minimal setup, we are able to design simple configurations of the weight matrices of a Transformer that enable it to solve fundamental problems such as electric flow and Laplacian eigenvector decomposition. Furthermore, we provide explicit error bounds that scale well with the number of Transformer layers. Several of our constructions rely crucially on the structure of the linear attention module, and may help shed light on the success of attention-based GNNs. We hope that our analysis paves the way to a better understanding of the learning landscape of graph Transformers, such as concrete bounds on their generalization and optimization errors.
Besides enhancing our understanding of Transformers, our results are also useful for the practical design of graph Transformers. In Sections 3 and 4, we show that essentially the same linear Transformer architecture is capable of learning a number of popular positional encodings (PE). In Section 6, we provide experimental evidence that the linear Transformer can learn better PEs than hard-coded PEs.
1.1 Summary of Contributions
Below, we summarize the main content of our paper.
  1. 1.
    We provide explicit weight configurations for which the Transformer implements efficient algorithms for several fundamental graph problems. These problems serve as important primitives in various graph learning algorithms, and have also been useful as PEs in state-of-the-art GNNs.
    1. (a)
      Lemma 1 constructs a Transformer that solves electric flows by implementing steps of gradient descent to minimize flow energy; consequently, it can also invert the graph Laplacian.
    2. (b)
      Lemmas 2 and 3 construct Transformers that compute low-dimensional resistive embeddings and heat kernels. Both constructions are based on implementing suitable power series.
    3. (c)
      By implementing a multiplicative polynomial expansion, Lemma 4 provides a construction for electric flow with exponentially higher accuracy than Lemma 1. Similarly, Lemma 5 provides a construction that computes the heat kernel using much fewer layers than Lemma 3.
    4. (d)
      In Lemma 6, we show that the Transformer can implement subspace iteration for finding the top-$k$ (or bottom-$k$) eigenvectors of the graph Laplacian. Central to this analysis is the ability of self-attention to compute a QR decomposition of the feature vectors.
    We derive explicit error bounds for the Transformer based on the convergence rates of the underlying algorithm implemented by the Transformer. We summarize these results in Table 1.
  2. 2.
    In Section 5, we provide a more efficient, constrained version of the linear Transformer with much fewer parameters. Although the constrained Transformer is less expressive by definition, we show that it can nonetheless implement all the above-mentioned constructions. Further, we show in Lemma 10 that the parameter-efficient linear Transformer has desirable invariance and equivariance properties.
  3. 3.
    We test the empirical performance of our theory on synthetic random graphs. In Section 3.5, we verify that Transformers with a few layers can achieve high accuracy in computing electric flows, resistive embeddings, as well as heat kernels. In Section 4.1, we verify that the Transformer can accurately compute top-$k$ and bottom-$k$ eigenevectors of the graph Laplacian.
  4. 4.
    In Section 6, we demonstrate the advantage of using the linear Transformer as a replacement for Laplacian eigenvector positional encoding on a molecular regression benchmark using the QM9 and ZINC datasets (Ruddigkeit et al., 2012; Ramakrishnan et al., 2014; Irwin et al., 2012). After replacing the Laplacian eigenvector-based PE with the linear Transformer, and training on the regression loss, we verify that the linear Transformer automatically learns a good PE for the downstream regression task that can outperform the original Laplacian eigenvector-based PE by a wide margin.
Lemma Task Transformer Implements: # layers for $\varepsilon$ error
1 Electric Flow (${\mathcal{L}}^{\dagger}$) Gradient Descent $\log(1/\varepsilon)$
4 Electric Flow (${\mathcal{L}}^{\dagger}$) Multiplicative Expansion $\log\log(1/\varepsilon)$
2 Resistive Embedding ($\sqrt{{\mathcal{L}}^{\dagger}}$) Power Series $\log(1/\varepsilon)$
3 Heat Kernel ($e^{-s{\mathcal{L}}}$) Power Series $\log(1/\varepsilon)+s\lambda_{\max}$
5 Heat Kernel ($e^{-s{\mathcal{L}}}$) Multiplicative Expansion $\log(1/\varepsilon)+\log(s\lambda_{\max})$
6 & 7 $k$-Eigenvector Decomp. of ${\mathcal{L}}$ Subspace Iteration (subspace-iteration steps)/$k$
Table 1: Summary of Error Bounds proved for various Transformer constructions. ${\mathcal{L}}$ is the graph Laplacian, ${\mathcal{L}}^{\dagger}$ denotes its pseudoinverse.
1.2 Related Work
Numerous authors have proposed different ways of adapting Transformers to graphs (Dwivedi and Bresson, 2020; Ying et al., 2021; Rampasek et al., 2022; Muller et al., 2023). A particularly promising approach is to use a positional encodings to incorporate structural information in the input, examples include Laplacian eigenvectors (Dwivedi and Bresson, 2020; Kreuzer et al., 2021), heat kernel (Choromanski et al., 2022), resistance distance and commute time (Ma et al., 2023; Velingker et al., 2024; Zhang et al., 2023) and shortest path distance (Ying et al., 2021). Lim et al. (2022) design a neural network to transform eigenvectors into an encoding that has certain invariance properties. Black et al. (2024) compare the expressivity of different PE schemes.
Kim et al. (2021) explore the possibility of using pure Transformers for graph learning. They provide both nodes and edges as input tokens, with a simple encoding scheme. They also prove that such a Transformer is as expressive as a second-order invariant graph network.
Ma et al. (2023) propose a novel graph Transformer architecture that incorporates learned PEs based on learned relative random walk probabilities (RRWP). Kreuzer et al. (2021) also propose learning PEs parameterized by eigenfunctions. In comparison, using a linear-Transformer to learn PEs has a few advantages: First, self attention can implement operations such as orthogonalization that are necessary for learning eigenvectors (Lemma 6). Second, self-attention enables some highly-efficient algorithms based on multiplicative polynomial expansions (Lemma 4). For these reasons, linear Transformers can learn good PEs given only the incidence matrix as input.
Finally, we note several other relevant papers that are unrelated to graph neural networks. A recent body of work proves the ability of Transformers to implement learning algorithms, to explain the in-context learning phenomenon (Schlag et al., 2021; Von Oswald et al., 2023). Surprisingly, our construction for Lemma 1 bears several remarkable parallels to the gradient descent construction by Von Oswald et al. (2023). We conjecture that the proof of Lemma 4 may be applicable for understanding the GD++ algorithm in the same paper. In another direction, Charton (2021) show experimentally that Transformers can compute linear algebra operations such as eigenvector decomposition. Their work requires a relatively complicated matrix encoding and a large number of parameters.
2 Preliminaries and Notation
2.1 Graphs
We use ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ to denote a graph with vertex set ${\mathcal{V}}$ and edge set ${\mathcal{E}}$; $n$ denotes $|{\mathcal{V}}|$ and $d$ represents $|{\mathcal{E}}|$. We often identify the vertex $v_{i}$ with its index $i$ for $i=1...n$, and the edge $e_{j}$ with $j$ for $j=1...d$. We will generally consider weighted graphs, where $r(\cdot):{\mathcal{E}}\to\mathbb{R}^{+}$ gives the edge weights. We will use $r_{j}:=r(e_{j})$. A central object of interest is the incidence matrix $B\in\mathbb{R}^{n\times d}$, defined as follows: to each edge $e_{j}\in{\mathcal{E}}$, assign an arbitrary orientation, i.e. $e=(u\to v)\in{\mathcal{E}}$. The matrix $B$ is given by
$B_{ij}=\left\{\begin{array}[]{ll}-1/\sqrt{r_{j}},&\text{ if exists $v\in{\mathcal{V}}$ such that $e_{j}=(u_{i}\to v)$}\\ +1/\sqrt{r_{j}},&\text{ if exists $v\in{\mathcal{V}}$ such that $e_{j}=(v\to u_{i})$}\\ 0,&\text{ otherwise.}\end{array}\right.$
Next, we define the graph Laplacian as ${\mathcal{L}}:=BB^{\top}\in\mathbb{R}^{n\times n}$. We will often need to refer to the maximum eigenvalue of ${\mathcal{L}}$, which we denote as $\lambda_{\max}$. Note that ${\mathcal{L}}$ always has 00 as its smallest eigenvalue, with corresponding eigenvector $\frac{1}{\sqrt{n}}\vec{1}$, the all-ones vector. This fact can be verified by noticing that each column of $B$ sums to 00. For a connected graph (as we will assume is the case throughout the paper), the second-smallest eigenvalue is always non-zero, and we will denote it as $\lambda_{\min}$.
Finally, we will frequently refer to the matrix $I_{n\times n}-\frac{1}{n}\vec{1}\vec{1}^{\top}$; for ease of reference, we will write $\hat{I}:=I_{n\times n}-\frac{1}{n}\vec{1}\vec{1}^{\top}$. This matrix is the projection onto $({\mathcal{L}})$, and it essentially functions as the identity matrix, as we work within $({\mathcal{L}})$ most of the time.
2.2 Linear Transformer
We will use $Z_{0}\in\mathbb{R}^{h\times n}$ to denote the input to the Transformer. $Z_{0}$ encodes a graph ${\mathcal{G}}$, and each column of $Z_{0}$ encodes a single vertex in $h$ dimensions. Let $W^{Q},W^{K},W^{V}\in\mathbb{R}^{h\times h}$ denote the key, query and value parameter matrices. We define linear attention $\mathrm{Attn}$ as
$\mathrm{Attn}_{W^{V},W^{Q},W^{K}}(Z):=W^{V}ZZ^{\top}{W^{Q}}^{\top}W^{K}Z.$ (2.2)
Unlike standard attention, (2.2) does not contain softmax activation. We construct an $L$-layer Transformer by stacking $L$ layers of the attention module (with linear feed-forward). To be precise, let $Z_{l}$ denote the output of the $(l-1)^{th}$ layer of the Transformer. Then
$Z_{l+1}:=Z_{l}+\mathrm{Attn}_{W^{V}_{l},W^{Q}_{l},W^{K}_{l}}{\left(Z_{l}\right)}+{W^{R}}Z_{l},$ (2.3)
where $W^{V}_{l},W^{Q}_{l},W^{K}_{l}$ are the value, query and key weight matrices of the linear attention module at layer $l$, and $W^{R}\in\mathbb{R}^{h\times h}$ is the weight matrix of the linear module. Henceforth, we let $W^{V}:=\{W^{V}_{l}\}_{l=0\ldots L}$, $W^{Q}:=\{W^{Q}_{l}\}_{l=0\ldots L}$, $W^{K}_{l}:=\{W^{K}_{l}\}_{l=0\ldots L}$, $W^{R}_{l}:=\{W^{R}_{l}\}_{l=0\ldots L}$ denote collections of the parameters across all layers of an $L$-layer Transformer.
Finally, we will often need to refer to specific rows of $Z_{l}$. We will use ${\left[Z_{l}\right]}_{i...j}$ to denote rows $i$ to $j$ of $Z_{l}$.
3 Transformers as powerful solvers for Laplacian problems
In this section, we discuss the capacity of the linear Transformer 2.3 to solve certain classes of canonical graph problems. We begin with the problem of Electric Flow in Section 3.2, where the constructed Transformer can be interpreted as implementing steps of gradient descent with respect to the energy of the induced flow. Subsequently, in Section 3.3, we provide constructions for computing the resistive embedding, as well as the heat kernel, based on implementing a truncated power series. Finally, in Section 3.4, we provide faster alternative constructions for solving electric flow and computing heat kernels, based on implementing a multiplicative polynomial expansion. In each case, we bound the error of the Transformer by the convergence rate of the underlying algorithms.
3.1 Additional Setup
We introduce some additional setup that is common to many lemmas in this section. We will generally consider an $L$-layer Transformer, as defined in (2.3), for some arbitrary $L\in\mathbb{Z}^{+}$. As in (2.3), we use $Z_{l}$ to denote the input to layer $l$. The input $Z_{0}$ will encode information about a graph ${\mathcal{G}}$, along with a number of demand vectors $\psi_{1}\ldots\psi_{k}\in\mathbb{R}^{n}$. We use $\Psi$ to denote the $n\times k$ matrix whose $i^{th}$ column is $\psi_{i}$. Unless otherwise stated, $Z_{l}\in\mathbb{R}^{(d+2k)\times n}$, where $n$ is the number of nodes, $d$ is the number of edges, and $k$ is the number of demands/queries. $Z_{0}^{\top}:=\begin{bmatrix}B,&\Psi,&0_{n\times k}\end{bmatrix}$.
On parameter size: In a straightforward implementation, the above Transformer has feature dimension $h=(d+2k)$. The size of $W^{Q},W^{K},W^{V},W^{R}$ is $O(h^{2})=O(d^{2}+k^{2})$, which is prohibitively large as $d$ can itself be $O(n^{2})$. The size of parameter matrices can be significantly reduced to $O(k^{2})$ by imposing certain constraints on the parameter matrices; we present this reduction in (5.1) in Section 5. For simplicity of exposition, lemmas in this section will use the naive implementation in (2.3). We verify later that all the constructions presented in this section can also be realized in (5.1).
3.2 Solving Electric Flow with Gradient Descent
Assume we are given a graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$, along with a non-negative vector of resistances $r\in\mathbb{R}_{+}^{d}$. Let $R$ be the $d\times d$ diagonal matrix with $r$ on its diagonal. A flow is represented by $f\in\mathbb{R}^{d}$, where $f_{j}$ denotes the (directed) flow on edge $e_{j}$. The energy of an electric flow is given by $\sum_{j=1}^{d}r_{j}f_{j}^{2}$. Let $\psi\in\mathbb{R}^{n}$ denote a vector of demands. Throughout this paper, we will assume that the demand vectors satisfy flow conservation, i.e., $\langle{\psi,\vec{1}}\rangle=0$. The $\psi$-electric flow is the unique minimizer of the following (primal) flow-optimization problem (by convex duality, this is equivalent to a dual potential-optimization problem):
(primal) $\min_{f\in\mathbb{R}^{d}}\quad\sum_{j=1}^{d}r_{j}f_{j}^{2}\qquad\qquad\text{subject to the constraint}\quad BR^{1/2}f=\psi.$ (3.1)
(dual) $-\min_{\phi\in\mathbb{R}^{n}}\phi^{\top}{\mathcal{L}}\phi-2\phi^{\top}\psi.$ (3.2)
The argument is standard; for completeness, we provide a proof of equivalence between (3.1) and (3.2) in (8.1) in Appendix 8.1. It follows that the optimizer $\phi^{*}$ of (3.2) has closed-form solution $\phi^{*}={\mathcal{L}}^{\dagger}\psi$. In Lemma 1 below, we show a simple construction that enables the Transformer in (2.3) to compute in parallel, the optimal potential assignments for a set of $k$ demands $\left\{\psi_{i}\right\}_{i=1...k}$, where $\psi_{i}\in\mathbb{R}^{n}$.
Motivation: Effective Resistance Metric
An important practical motivation for solving the electric flow (or equivalently computing ${\mathcal{L}}^{\dagger}$) is to obtain the Effective Resistance matrix ${\mathcal{R}}\in\mathbb{R}^{n\times n}$. GNNs that use positional encodings derived from ${\mathcal{R}}$ have demonstrated state-of-the-art performance on numerous tasks, and can be shown to have good theoretical expressivity (Zhang et al., 2023; Velingker et al., 2024; Black et al., 2024). Formally, ${\mathcal{R}}$ is defined as ${\mathcal{R}}_{ij}:=(u_{i}-u_{j})^{\top}{\mathcal{L}}^{\dagger}(u_{i}-u_{j})$, where $u_{i}$ denotes the vector that has a $1$ in the $i^{th}$ coordinate, and 00s everywhere else. Intuitively, ${\mathcal{R}}_{ij}$ is the potential drop required to send 1-unit of electric flow from node $i$ to node $j$. Let $\ell\in\mathbb{R}^{n}$ denote the vector of diagonals of ${\mathcal{L}}^{\dagger}$ (i.e. $\ell_{i}:={\mathcal{L}}_{ii}$). Then ${\mathcal{R}}=\vec{1}\ell^{\top}+\ell\vec{1}^{\top}-2{\mathcal{L}}^{\dagger}$; thus computing ${\mathcal{L}}^{\dagger}$ essentially also computes ${\mathcal{R}}$.
Lemma 1 (Transformer solves Electric Flow by implementing Gradient Descent).
Consider the setup in Section 3.1. Assume that $\langle{\psi_{i},\vec{1}}\rangle=0$ for each $i=1...k$. For any $\delta>0$ and for any $L$-layer Transformer, there exists a choice of weights $W^{V},W^{Q},W^{K},W^{R}$, such that each layer of the Transformer (2.3) implements a step of gradient descent with respect to the dual electric flow objective in (3.2), with stepsize $\delta$. Consequently, the following holds for all $i=1...k$ and for any graph Laplacian with maximum eigenvalue $\lambda_{\max}\leq 1/\delta$ and minimum nontrivial eigenvalue $\lambda_{\min}$:
${\left\|{\left[Z_{L}\right]}_{d+k+i}^{\top}-{\mathcal{L}}^{\dagger}\psi_{i}\right\|}_{2}\leq\frac{\exp{\left(-\delta\lambda_{min}/2\right)}}{\sqrt{\lambda_{min}}}\|\psi_{i}\|_{2}^{2}{\left\|\psi_{i}\right\|}_{2}.$
Discussion. In the special case when $k=n$ and $\Psi:=I_{n\times n}-\frac{1}{n}\vec{1}\vec{1}^{\top}$, ${\left[Z_{L}\right]}_{d+k+1...d+2k}\approx{\mathcal{L}}^{\dagger}$. An $\varepsilon$-approximate solution requires $L=O(\log(1/\varepsilon))$ layers; this is exactly the convergence rate of gradient descent on a Lipschitz smooth and strongly convex function. In Lemma 4 below, we show an alternate construction that reduces this to $O(\log\log(1/\varepsilon))$.
We defer details of the proof of Lemma 1 to Appendix 8.1, and we provide the explicit choice of weight matrices in (8.3). The constructed weight matrices are very sparse; each of $W^{V},W^{Q},W^{K},W^{R}$ contains a single identity matrix in a sub-block. This sparse structure makes it possible to drastically reduce the number of parameters needed, which we exploit in Section 5 to design a more parameter efficient Transformer. We provide experimental validation of Lemma 1 in Section 3.5.
3.3 Implementing Truncated Power Series
In this section, we present two more constructions: one for computing $\sqrt{{\mathcal{L}}^{\dagger}}$ (Lemma 2), and one for computing the heat kernel $e^{-s{\mathcal{L}}}$ (Lemma 3). Both quantities have been successfully used for positional encoding in various GNNs. The constructions proposed in these two lemmas are also similar, and involve implementing the power series of the respective targets.
3.3.1 Computing The Principal Square Root $\sqrt{{\mathcal{L}}^{\dagger}}$
Motivation : Resistive Embedding
The following fact relates the effective resistance matrix ${\mathcal{R}}$ to any “square root” of ${\mathcal{L}}^{\dagger}$:
Fact 1.
Let ${\mathcal{M}}$ denote any matrix that satisfies ${\mathcal{M}}{\mathcal{M}}^{\top}={\mathcal{L}}^{\dagger}$. Let +${\mathcal{R}}\in\mathbb{R}^{n\times n}_{+}$ denote the matrix of effective resistances (see Section 3.2). Then ${\mathcal{R}}_{ij}={\left\|{\mathcal{M}}_{i}-{\mathcal{M}}_{j}\right\|}_{2}^{2}$, where ${\mathcal{M}}_{i}$ is the $i^{th}$ row of ${\mathcal{M}}$.
One can verify the above by noticing that ${\left\|{\mathcal{M}}_{i}-{\mathcal{M}}_{j}\right\|}_{2}^{2}={\mathcal{M}}_{i}^{\top}{\mathcal{M}}_{i}+{\mathcal{M}}_{j}^{\top}{\mathcal{M}}_{j}-2{\mathcal{M}}_{i}^{\top}{\mathcal{M}}_{j}={\left[{\mathcal{L}}\right]}_{ii}+{\left[{\mathcal{L}}\right]}_{jj}-2{\mathcal{L}}_{ij}$. In some settings, it is more natural to use the rows of ${\mathcal{M}}$ to embed node position, instead of directly using ${\mathcal{R}}$: By assigning an embedding vector $w_{i}:={\mathcal{M}}_{i}$ to vertex $i$, the Euclidean distance between $w_{i}$ and $w_{j}$ equals the resistance distance.
The matrix ${\mathcal{M}}$ is under-determined, and for any $m>n$, there are infinitely many choices of ${\mathcal{M}}$ that satisfy ${\mathcal{M}}{\mathcal{M}}^{\top}={\mathcal{L}}^{\dagger}$. Fact 1 applies to any such ${\mathcal{M}}$. Velingker et al. (2024) uses the rows of ${\mathcal{M}}={\mathcal{L}}^{\dagger}B$ for resistive embedding. Under this choice, ${\mathcal{M}}_{i}$ has dimension $d$, which is the number of edges and can be quite large. To deal with this, Velingker et al. (2024) additionally performs a dimension-reduction step using Johnson Lidenstrauss.
Among all valid choices of ${\mathcal{M}}$, there is a unique choice that is symmetric and minimizes $\|{\mathcal{M}}\|_{F}$, namely, $US^{-1/2}U^{\top}$, where $USU^{\top}={\mathcal{L}}$ is the eigenvector decomposition of ${\mathcal{L}}$. We reserve $\sqrt{{\mathcal{L}}^{\dagger}}$ to denote this matrix; $\sqrt{{\mathcal{L}}^{\dagger}}$ is called the principal square root of ${\mathcal{L}}^{\dagger}$. In practice, $\sqrt{{\mathcal{L}}^{\dagger}}$ might be preferable to, say, ${\mathcal{L}}^{\dagger}B$ because it has an embedding dimension of $n$, as opposed to the possibly much larger $d$. We present in Lemma 2 a Transformer construction for computing $\sqrt{{\mathcal{L}}^{\dagger}}$.
Lemma 2 (Principal Square Root $\sqrt{{\mathcal{L}}^{\dagger}}$).
Consider the setup in Section 3.1. Assume that $\psi_{1}...\psi_{k}\in\mathbb{R}^{n}$ satisfy $\langle{\psi_{i},\vec{1}}\rangle=0$. For any $L$-layer Transformer (2.3), there exists a configuration of weights $W^{V},W^{Q},W^{K},W^{R}$ such that the following holds: For any graph with maximum Laplacian eigenvalue less than $\lambda_{\max}$ and minimum non-trivial Laplacian eigenvalue greater than $\lambda_{\min}$, and for all $i=1...k$:
${\left\|{\left[Z_{L}\right]}_{d+k+i}^{\top}-\sqrt{{\mathcal{L}}^{\dagger}}\psi_{i}\right\|}_{2}\leq\frac{e^{-L\lambda_{\min}/\lambda_{\max}}}{\lambda_{\min}\sqrt{L/\lambda_{\max}}}{\left\|\psi_{i}\right\|}_{2}.$
Discussion. We defer a proof of Lemma 2 to Appendix 8.2. The high-level idea is that each layer of the Transformer implements one additional term of the power series expansion of $\sqrt{{\mathcal{L}}^{\dagger}}$. Under the choice $k=n$ and $\Psi=I_{n\times n}-(1/n)\vec{1}\vec{1}^{\top}$, ${\left[Z_{L}\right]}_{d+k+1...d+2k}\approx\sqrt{{\mathcal{L}}^{\dagger}}$. An $\varepsilon$ approximation requires $\log(1/\varepsilon)$ layers. We consider the more general setup involving arbitrary $\psi_{i}^{\prime}s$ as they are useful for projecting the resistive embedding onto a lower-dimensional space; this is relevant when $\psi_{i}$’s are trainable parameters (see e.g., the setup in Appendix 12.3). We empirically validate Lemma 2 in Section 3.5.
3.3.2 Computing the heat kernel: $\exp{\left(-s{\mathcal{L}}\right)}$
Finally, we present a result on learning heat kernels. The heat kernel has connections to random walks and diffusion maps Coifman and Lafon (2006). It plays a central role in semi-supervised learning on graphs (Xu et al., 2020); it is also used for positional encoding (Choromanski et al., 2022). We note that in some settings, the heat kernel excludes the nullspace of ${\mathcal{L}}$ and is instead defined as $e^{-s{\mathcal{L}}}-\vec{1}\vec{1}^{\top}/n$.
Lemma 3 (Heat Kernel $e^{-s{\mathcal{L}}}$).
Consider the setup in Section 3.1. Assume that $\psi_{1}...\psi_{k}\in\mathbb{R}^{n}$ satisfy $\langle{\psi_{i},\vec{1}}\rangle=0$. Let $s>0$ be an arbitrary temperature parameter. There exists a configuration of weights for the $L$-layer Transformer (2.3) such that the following holds: for any input graph whose Laplacian ${\mathcal{L}}$ satisfies $8s\lambda_{\max}\leq L$, and for all $i=1...k$,
${\left\|{\left[Z\right]}_{L,i}^{\top}-e^{-s{\mathcal{L}}}\psi_{i}\right\|}_{2}\leq 2^{-L+8s\lambda_{\max}+1}{\left\|\psi_{i}\right\|}_{2}$
Discussion. We defer the proof of Lemma 3 to Appendix 8.3. To obtain $\varepsilon$ approximation error, we need number of layers $L\geq{O{\left(\log(1/\varepsilon)+s\lambda_{\max}\right)}}$. As with the preceding lemmas, the flexibility of choosing any number of $\psi_{i}$’s enables the Transformer to learn a low-dimensional projection of the heat kernel. The dependence on $s\lambda_{\max}$ is fundamental, stemming from the fact that the truncation error of the power series of $e^{-s{\mathcal{L}}}$ begins to shrink only after $O(s\lambda_{\max})$ the first terms. In Lemma 5 in the next section, we weaken this dependence from $s\lambda_{\max}$ to $\log(s\lambda_{\max})$. We empirically validate Lemma 3 in Section 3.5.
3.4 Implementing Multiplicative Polynomial Expansion
We present two alternative Transformer constructions that can achieve vastly higher accuracy than the ones in preceding sections: Lemma 4 computes an $\varepsilon$-accurate electric flow in exponentially fewer layers than Lemma 1. Lemma 5 approximates the heat kernel with higher accuracy than Lemma 3 when the number of layers is small. The key idea in both Lemmas 4 and 5 is to use the Transformer to implement a multiplicative polynomial; this in turn makes key use of the self-attention module.
The setup for Lemmas 4 and 5 differs in two ways from that presented in Section 3.1. First, the input to layer $l$, $Z_{l}$, are now in $\mathbb{R}^{3n\times n}$, instead of $\mathbb{R}^{(d+2k)\times n}$. When the graph ${\mathcal{G}}$ is sparse and the number of demands/queries $k$ is small, the constructions in Lemma 1 and 3 may use considerably less memory. This difference is fundamental, due to the storage required for raising matrix powers. Second, the input is also different; in particular, information about the graph is provided via ${\mathcal{L}}$ as part of the input $Z_{0}$, as opposed to via the incidence matrix $B^{\top}$ as done in Section 3.1. This difference is not fundamental: one can compute ${\mathcal{L}}=B^{\top}B$ from $B$ in a single attention layer; in the spirit of brevity, we omit this step. We begin with the faster construction for electric flow:
Lemma 4.
Let $\delta>0$. Let $Z_{0}^{\top}:=\begin{bmatrix}(\hat{I}_{n\times n}-\delta{\mathcal{L}}),&I_{n\times n},&\delta I_{n\times n}\end{bmatrix}$. Then there exist a choice of $W^{V},W^{Q},W^{K},W^{R}$ for a $L$-layer Transformer (2.3) such that for any graph with Laplacian with smallest non-trivial eigenvalue $\lambda_{\min}$ and largest eigenvalue $\lambda_{\max}\leq 1/\delta$,
${\left\|{\left[Z_{L}\right]}_{2n+1...3n}-{\mathcal{L}}^{\dagger}\right\|}_{2}\leq\frac{1}{\lambda_{\min}}\exp{\left(-\delta 2^{L}\lambda_{\min}\right)}.$
Discussion. Lemma 4 shows that the Transformer can compute an $\varepsilon$-approximation to ${\mathcal{L}}^{\dagger}$ (which is sufficient but not necessary for solving arbitrary electric flow demands) using $\log\log(1/\varepsilon)$ layers. This is much fewer than the $\log(1/\varepsilon)$ layers required in Lemma 1. The key idea in the proof is to implement a multiplicative polynomial expansion for ${\mathcal{L}}^{\dagger}$. We defer the proof to Appendix 8.4.
Next, we show the alternate construction for computing the heat kernel:
Lemma 5 (Fast Heat Kernel).
Let $s>0$. Let $L$ be the number of Transformer layers. Let $Z_{0}^{\top}:=(I_{n\times n}-3^{-L}s{\mathcal{L}})$. Then there exist a choice of $W^{V},W^{Q},W^{K},W^{R}$ such that for any graph whose Laplacian ${\mathcal{L}}$ satisfies $s\lambda_{\max}\leq 3^{L}$,
${\left\|Z_{L}-\exp(-s{\mathcal{L}})\right\|}_{2}\leq 3^{-L+1}{s^{2}\lambda_{\max}^{2}}.$
Discussion. Lemma 5 shows that the Transformer can compute an $\varepsilon$-approximation to $e^{-s{\mathcal{L}}}$ using $O(\log(1/\varepsilon)+\log(s\lambda_{\max}))$ layers. The $\varepsilon$ dependence is the same as Lemma 3, but the $\log(s\lambda_{\max})$ is an improvement. When the number of layers is small, Lemma 5 gives a significantly more accurate approximation. The proof is based on the well-known approximation $e^{x}\approx(1+x/L)^{L}$. We defer proof details for Lemma 5 to Appendix 8.4.
3.5 Experiments
In Figure 1, we experimentally verify that the Transformer is capable of learning to solve the three objectives presented in Lemmas 1, 2 and 3. The setup is as described in Section 3.1, with the Transformer described in (2.3) with $k=n$, but with the following important difference: the output of each layer is additionally normalized as follows: ${\left[Z_{l}\right]}_{1...n}\leftarrow{\left[Z_{l}\right]}_{1...n}/\|{\left[Z_{l}\right]}_{1...n}\|_{F}$, ${\left[Z_{l}\right]}_{n+1...n+2k}\leftarrow{\left[Z_{l}\right]}_{n+1...n+2k}/\|{\left[Z_{l}\right]}_{n+1...n+2k}\|_{F}$, where $\|\|_{F}$ is the Frobenius norm. Without this normalization, training the Transformer becomes very difficult beyond 5 layers.
We consider two kinds of random graphs: fully-connected graphs ($n=10,d=45$) and Circular Skip Links (CSL) graphs ($n=10,d=20$). Edge resistances are randomly sampled. We provide details on the sampling distributions in Appendix 11. For each input graph ${\mathcal{G}}$, we sample $n$ demands $\psi_{1}...\psi_{n}\in\mathbb{R}^{n}$ independently from the unit sphere. Let $\Psi={\left[\psi_{1}...\psi_{n}\right]}$. The input to the Transformer is $\begin{bmatrix}B^{\top},&\Psi^{\top},&0_{n\times n}\end{bmatrix}$, consistent with the setup described in Section 3.1. The training/test loss is given by $\text{loss}_{U}:=\mathbb{E}{\left[\frac{1}{n}\sum_{i=1}^{n}{\left\|\frac{{\left[Z_{l}\right]}_{d+n+i}^{\top}}{\|{\left[Z_{l}\right]}_{d+n+i}\|_{2}}-\frac{U\psi_{i}}{\|U\psi_{i}\|_{2}}\right\|}_{2}^{2}\right]},$ where $U\in\left\{{\mathcal{L}}^{\dagger},\sqrt{{\mathcal{L}}^{\dagger}},e^{-0.5{\mathcal{L}}}\right\}$. We learn the correct solutions only up to scaling, because we need to normalize $Z_{l}$ per-layer. The expectation is over randomness in sampling ${\mathcal{L}}$ and $\Psi$. We plot the log of the respective losses against number of layers after training has converged. As can be seen, in each plot, and for both types of architectures, the loss appears to decrease exponentially with the number of layers.
Refer to caption Refer to caption Refer to caption
1(a) $\text{loss}_{{\mathcal{L}}^{\dagger}}$ (electric) 1(b) $loss_{\sqrt{{\mathcal{L}}^{\dagger}}}$ (resistive) 1(c) $\text{loss}_{\exp{\left(-0.5{\mathcal{L}}\right)}}$ (heat)
Figure 1: $\text{loss}_{U}$ against number of layers at convergence for $U\in\left\{{\mathcal{L}}^{\dagger},\sqrt{{\mathcal{L}}^{\dagger}},e^{-0.5{\mathcal{L}}}\right\}$.
4 Transformers can Implement Subspace Iteration to Compute Eigenvectors
We present a Transformer construction for finding the eigenvectors of the graph Laplacian ${\mathcal{L}}$. The eigenvector for the smallest non-trivial eigenvalue, for example, has been applied with great success in graph segmentation (Shi and Malik, 1997) and clustering (Bühler and Hein, 2009). Additionally, the Laplacian eigenvectors are also commonly used for positional encoding (see Section 1.2). Our construction is based on the subspace iteration algorithm (Algorithm 1), aka block power method—see e.g., (Bentbib and Kanber, 2015). The output $\Phi$ of Algorithm 1 converges to the top-$k$ eigenvectors of ${\mathcal{L}}$. We show in Corollary 7 that a modified construction can also find the bottom-$k$ eigenvectors of ${\mathcal{L}}$.
Algorithm 1 Subspace Iteration
$\Phi_{0}\in\mathbb{R}^{n\times k}$ has full column rank .
while not converged do
     $\hat{\Phi}\leftarrow{\mathcal{L}}\Phi$
     $\Phi\leftarrow QR(\Phi)$triangleright QR($\Phi$) returns the Q matrix in the QR decomposition of $\Phi$.
end while
For the purposes of this section, we consider a variant of the Transformer defined in (2.3).
$\hat{Z}_{l+1}:=Z_{l}+\mathrm{Attn}_{W^{V}_{l},W^{Q}_{l},W^{K}_{l}}{\left(Z_{l}\right)}+{W^{R}}Z_{l}$
${Z}_{l+1}=\text{normalize}(\hat{Z}_{l+1}),$ (4.1)
where $\text{normalize}(Z)$ applies row-wise normalization: ${\left[\text{normalize}(Z)\right]}_{i}\leftarrow{\left[Z\right]}_{i}/\|{\left[Z\right]}_{i}\|_{2}$ for $i=d+1...d+k$. We use $B_{l}\in\mathbb{R}^{n\times d}$ and $\Phi_{l}\in\mathbb{R}^{n\times k}$ to denote refer to specific columns of $Z_{l}^{\top}$, defined as $Z_{l}^{\top}=:\begin{bmatrix}B_{l}^{T},&\Phi_{l}^{\top}\end{bmatrix}$. The notation is chosen as $\Phi_{l}$ corresponds to $\Phi$ in Algorithm 1. We initialize $B_{0}=B$ and let $\Phi_{0}$ be any column-orthogonal matrix.
Lemma 6 (Subspace Iteration for Finding Top $k$ Eigenvectors).
Consider the Transformer defined in (4.1). There exists a choice of $W^{V},W^{Q},W^{K},W^{R}$ such that $k+1$ layers of the Transformer implements one iteration of Algorithm 1. Consequently, the output $\Phi_{L}$ of a $L$-layer Transformer approximates the top-$k$ eigenvectors of ${\mathcal{L}}$ to the same accuracy as $L/(k+1)$ steps of Algorithm 1.
Discussion. We bound the Transformer’s error by the error of Algorithm 1, but do not provide an explicit convergence rate. This omission is because the convergence rate of Algorithm 1 itself is difficult to characterize, and has a complicated dependence on pairwise spectral gaps. The high-level proof idea is to use self-attention to orthogonalize the columns of $\Phi_{l}$ (see Lemma 8). We defer the proof of Lemma 6 to Appendix 9. We experimentally validate Lemma 6 in Section 4.1.
The construction in Lemma 6 explicitly requires $k$ layers to perform a QR decomposition of $\Phi$ plus 1 layer to multiply by ${\mathcal{L}}$. The layer usage can be much more efficient in practice: First, multiple ${\mathcal{L}}$-multiplications can take place before a single QR-factorization step. Second, the $k$ layers for performing QR decomposition can be implemented in a single layer with $k$ parallel heads.
In graph applications, one is often interested in the smallest-$k$ eigenvectors of ${\mathcal{L}}$:
Corollary 7 (Subspace Iteration for Finding Bottom $k$ Eigenvectors).
Consider the same setup as Lemma 6. Let $\mu>0$ be some constant. There exists a construction for a $L$-layer Transformer (similar to Lemma 6), which implements Algorithm 1 with ${\mathcal{L}}$ replaced by $\mu I-{\mathcal{L}}$, and $\Phi_{L}$ approximates the smallest $k$ eigenvectors of ${\mathcal{L}}$ if $\lambda_{\max}({\mathcal{L}})\leq\mu$.
We provide a short proof in Appendix 9. The key idea is that a minor modification of the contruction in Lemma 6 will instead compzute the bottom $k$ eigenvectors. Alternative to the construction in Corollary 7, one can also first compute ${\mathcal{L}}^{\dagger}$ (via Lemma 4), and then perform subspace iteration for ${\mathcal{L}}^{\dagger}$.
4.1 Experiments for Lemma 6
We verify Lemma 6 and Corollary 7 experimentally by evaluating the ability of the Transformer (2.3) to learn top-$k$ and bottom-$k$ eigenvectors. As in Section 3.5, we consider two kinds of random graphs with $n=10$ nodes: fully connected ($d=45$ edges) and CSL ($d=20$ edges); each edge is has a randomly sampled resistance; see Appendix 11 for details. For a graph ${\mathcal{G}}$ with Laplacian ${\mathcal{L}}$, let $\lambda_{1}\leq\lambda_{2}\leq...\lambda_{10}$ denote its eigenvalues. Let $v_{1},\dots,v_{10}$ denote its eigenvectors. $\lambda_{1}$ is always 00 and $v_{1}$ is always $\vec{1}/\sqrt{n}$.
The Transformer architecture is as defined in Section 4.1, with $k=n$. We increase the dimension of $\Phi_{l}$ to $(2n)\times n$, and thus the dimension of $Z_{l}$ to $(d+n+n)\times n$. We read out the last $n$ rows of $Z_{L}$ as output. The purpose of increasing the dimension is to make the architecture identical to the one used in the experiments in Section 3.5; the construction in Lemmas 6 and Corollary 7 extend to this setting by setting appropriate parameters to 00. In addition, we also normalize ${\left[Z_{l}\right]}_{1...d}$ each layer by its Frobenius norm, i.e., ${\left[Z_{l}\right]}_{1...d}\leftarrow{\left[Z_{l}\right]}_{1...d}/\|{\left[Z_{l}\right]}_{1...d}\|_{F}$ (the proof of Lemma 6 still holds as subspace iteration is scaling-invariant). The input to the Transformer is $Z_{0}^{\top}=\begin{bmatrix}B,&\Phi_{0}\end{bmatrix}$, and we make $\Phi_{0}\in\mathbb{R}^{2n\times n}$ a trainable parameter along with $W^{V},W^{Q},W^{K},W^{R}$. We define $\text{loss}_{i}:=\mathbb{E}{\left[\min\left\{\|\phi_{L,i}-v_{i}\|_{2}^{2},\|\phi_{L,i}+v_{i}\|_{2}^{2}\right\}\right]}$, where $\phi_{L,i}$ is the $i^{th}$ row of $\Phi_{L}$ and expectation is taken with respect to randomness in sampling the graph; the $\min$ is required because eigenvectors are invariant to sign-flips. We train and evaluate the Transformer on two kinds of losses: $\text{loss}_{1-5}:=\frac{1}{5}\sum_{i=1}^{5}\text{loss}_{i}$ and $\text{loss}_{1-10}:=\frac{1}{10}\sum_{i=1}^{10}\text{loss}_{i}$. We plot the results in Figure 2, and summarize our findings below:
  1. 1.
    2(a) and 2(d): Both $\text{loss}_{1-5}$ and $\text{loss}_{1-10}$ appear to decrease exponentially with the number of Transformer layers, for both FC and CSL graphs. This is consistent with Lemma 6 and Corollary 7.
  2. 2.
    In Figures 2(e) and 2(f), when the Transformer is trained on $\text{loss}_{1-10}$, larger eigenvectors are learned more accurately, with the exception of $loss_{2}$. This is mostly consistent with our construction in Lemma 6, where the larger eigenvectors are computed more quickly.
    In contrast, in Figures 2(b) and 2(c), when the Transformer is trained on $\text{loss}_{1-5}$, the smaller eigenvectors are learned more accurately. This is consistent with our construction in Corollary 7, where the smaller eigenvectors of ${\mathcal{L}}$ are also the larger eigenvectors of $\alpha I-{\mathcal{L}}$, and are thus computed more quickly. One expects to observe the same ordering if the Transformer is instead computing ${\mathcal{L}}^{\dagger}$ first, followed by subspace iteration.
We omit $\text{loss}_{1}$ from Figures 2(b), 2(c), 2(e), 2(f) because $v_{1}$ is a constant so $\text{loss}_{1}$ goes to 00 extremely fast, making it hard to see the other lines.
Refer to caption Refer to caption Refer to caption
2(a) $\text{loss}_{1-5}$    2(b) $\left\{\text{loss}_{i}\right\}_{i=2...5}$, FC 2(c) $\left\{\text{loss}_{i}\right\}_{i=2...5}$, CSL
Refer to caption Refer to caption Refer to caption
2(d) $\text{loss}_{1-10}$ 2(e) $\left\{\text{loss}_{i}\right\}_{i=2...10}$, FC 2(f) $\left\{\text{loss}_{i}\right\}_{i=2...10}$, CSL
Figure 2: $\log(\text{loss}_{*})$ vs. number of layers. Top row: various losses for the Transformer trained on $\text{loss}_{1-5}$. Bottom row: various losses for the Transformer trained on $\text{loss}_{1-10}$.
5 Parameter-Efficient Implementation
As explained in Section 3.1, the sizes of the $W^{Q},W^{K},W^{V},W^{R}$ matrices can scale with $O(n^{4})$ in the worst case, where $n$ is the number of nodes. This makes the memory requirement prohibitive. To alleviate this, we present below in (5.1) a more efficient Transformer implementation than the one defined in (2.3). (5.1) is strictly more constrained than (2.3), but it is still expressive enough to implement all the previous constructions. For each layer $l$, let $\alpha^{V}_{l}$, $\alpha^{Q}_{l}$, $\alpha^{K}_{l}$, $\alpha^{R}_{l}$ be scalar weight parameters, and let $W_{l}^{V,\Phi}$, $W_{l}^{Q,\Phi}$, $W_{l}^{K,\Phi}$, $W_{l}^{R,\Phi}\in\mathbb{R}^{2k\times 2k}$. Let $B_{l},\Phi_{l}$ evolve as
$B_{l+1}^{\top}=(1+\alpha^{R}_{l})B_{l}^{\top}+\alpha^{V}_{l}B_{l}^{\top}{\left(\alpha^{Q}_{l}\alpha^{K}_{l}B_{l}B_{l}^{\top}+\Phi_{l}{W_{l}^{Q,\Phi}}^{\top}W_{l}^{K,\Phi}\Phi_{l}^{\top}\right)}$ (5.1)
$\Phi_{l+1}^{\top}={\left(I+W_{l}^{R,\Phi}\right)}\Phi_{l}^{\top}+W_{l}^{V,\Phi}\Phi_{l}^{\top}{\left(\alpha^{Q}_{l}\alpha^{K}_{l}B_{l}B_{l}^{\top}+\Phi_{l}{W_{l}^{Q,\Phi}}^{\top}W_{l}^{K,\Phi}\Phi_{l}^{\top}\right)},$
with initialization $B_{0}:=B$, and $\Phi_{0}=[\Psi,0_{k\times n}]$. We verify that (5.1) matches (2.3), with $Z_{l}=\begin{bmatrix}B_{l}^{\top},\Phi_{l}^{\top}\end{bmatrix}$; the difference is that certain blocks of $W^{V}_{l}$, $W^{Q}_{l}$, $W^{K}_{l}$, $W^{R}_{l}$ constrained to be zero or some scaling of identity. Nonetheless, we verify that the constructions in Lemmas 1, 2, 3 and 6 can all be realized within the constrained dynamics (5.1). We prove this in Lemma 9 in Appendix 10.
Besides the reduced parameter size, we highlight two additional advantages of (5.1):
  1. 1.
    Let ${\sf TF}^{B}_{l}({\color[rgb]{0,0,1}B},{\color[rgb]{0,0,1}\Phi})$ (resp ${\sf TF}^{\Phi}_{l}({\color[rgb]{0,0,1}B},{\color[rgb]{0,0,1}\Phi}))$ be defined as the value of $B_{l}^{\top}$ (resp $\Phi_{l}^{\top}$) when initialized at $B_{0}={\color[rgb]{0,0,1}B}$ and $\Phi_{0}={\color[rgb]{0,0,1}\Phi}$ under the dynamics (5.1). Let $U\in\mathbb{R}^{d\times d}$ be some permutation matrix over edge indices. Then
    1. (a)
      ${\sf TF}^{B}_{l}$ is equivariant to edge permutation, i.e. ${\sf TF}^{B}_{l}(BU,\Phi)=U^{\top}{\sf TF}^{B}_{l}(B,\Phi)$
    2. (b)
      ${\sf TF}^{\Phi}_{l}$ is invariant to edge permutation, i.e. ${\sf TF}^{\Phi}_{l}(BU,\Phi)={\sf TF}^{\Phi}_{l}(B,\Phi)$.
    We provide a short proof of this in Lemma 10 in Appendix 10.
  2. 2.
    If $B_{l}$ is sparse, then $B_{l}^{\top}B_{l}$ can be efficiently computed using sparse matrix multiply. This is the case for all of our constructions for all layers $l$.
6 Learning Positional Encoding for Molecular Regression
In Sections 3.5 and 4.1, we saw that the Transformer is capable of learning a variety of embeddings based on ${\mathcal{L}}^{\dagger},\sqrt{{\mathcal{L}}},e^{-s{\mathcal{L}}}$, and ${\mathcal{L}}$), when the training loss is explicitly the recovery loss for that embedding. An natural question is then: Can a GNN perform as well on a downstream task if we replace its PE by a linear Transformer, and train the GNN together with the linear Transformer? There are two potential benefits to this approach: First, the Transformer can learn a PE that is better-suited to the task than hard-coded PEs, and thus achieve higher accuracy. Second, the Transformer with a few layers/dimensions may learn to compute only the most relevant features, and achieve comparable accuracy to the hard-coded PE using less computation.
To test this, we evaluate the performance of our proposed linear Transformer on a molecular regression task on two real-world datasets: QM9 (Ruddigkeit et al., 2012; Ramakrishnan et al., 2014) and ZINC (Irwin et al., 2012). The regression target is constrained solubility (Dwivedi et al., 2023). Our experiments are based on the Graph Transformer (GT) implementation from Dwivedi and Bresson (2020). In Table 2, we compare three loss values: GT without PE, GT with Laplacian Eigenvector as PE (LapPE), and GT with Linear Transformer output as PE. We choose to use LapPE as a baseline as it was the original PE of choice in Dwivedi and Bresson (2020). We present details of the datasets in Appendix 12.1. The architecture of linear Transformer is a modified version of (5.1), detailed in Appendix 12.2. Other experiment details, including precise definitions of each model, can be found in Appendix 12.3.
Model # Parameters Loss
ZINC Graph Transformer $800771-891$ $0.286\pm 0.0078$
Graph Transformer + LapPE $800771$ $0.201\pm 0.0034$
Graph Transformer + Linear Transformer $800771+488$ $0.138\pm 0.012$
QM9 Graph Transformer $799747-512$ $0.419\pm 0.0047$
Graph Transformer + LapPE $799747$ $0.227\pm 0.0094$
Graph Transformer + Linear Transformer $799747+240$ $0.221\pm 0.0060$
Table 2: Regression Loss for ZINC and QM9 for different choices of PE.
The difference between GT and GT + LapPE is substantial for both QM9 and ZINC, highlighting the importance of PE in both cases.
Going from GT + LapPE to GT + Linear Transformer for ZINC, there is a significant further improvement in loss (a drop of 0.063). This is remarkable, considering that the linear Transformer accounts for less than $0.7\%$ of the total number of parameters. Note that the SOTA error for ZINC regression is significantly lower than 0.138 on more recent architectures (see e.g. Lim et al. (2022)); the significance of our result is in demonstrating the improvement just by replacing the LapPE with the linear Transformer, while keeping everything else fixed.
In contrast, when going from GT + LapPE to GT + Linear Transformer for QM9, the difference is essentially zero. We conjecture that this may be because QM9 molecules (average of about $9$ nodes and $19$ edges) are considerably simpler than ZINC (average of about $23$ nodes and $40$ edges). Thus there may not be too many additional useful features to learn, and LapPE close to optimal. It is still consistent with our theory, for the Linear Transformer to do as well as LapPE.
7 Acknowledgement
The authors would like to thank Professor Xavier Bresson from the National University of Singapore for providing the GitHub code which facilitated some of the experiments in this work. Suvrit Sra acknowledges generous support from the Alexander von Humboldt Foundation.